OpenStudy (anonymous):

Which point lies on a circle that is centered at A(-3, 2) and passes through B(1, 3)?

OpenStudy (jdoe0001):

check the distance between the center and a point on the circle, thus getting the RADIUS so any point whose distance from the center equals the RADIUS, thus lies on the circle \(\bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ A&({\color{red}{ -3}}\quad ,&{\color{blue}{ 2}})\quad B&({\color{red}{ 1}}\quad ,&{\color{blue}{ 3}}) \end{array}\qquad d = \sqrt{({\color{red}{ x_2}}-{\color{red}{ x_1}})^2 + ({\color{blue}{ y_2}}-{\color{blue}{ y_1}})^2} \)

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

so C -1,-2

OpenStudy (jdoe0001):

hmmm what did you get for the radius?

OpenStudy (anonymous):

r=5

OpenStudy (jdoe0001):

\(\bf \bf \textit{distance between 2 points}\\ \quad \\ \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ A&({\color{red}{ -3}}\quad ,&{\color{blue}{ 2}})\quad B&({\color{red}{ 1}}\quad ,&{\color{blue}{ 3}}) \end{array}\qquad d = \sqrt{({\color{red}{ 1}}-{\color{red}{ -3}})^2 + ({\color{blue}{ 3}}-{\color{blue}{ 2}})^2}\)

OpenStudy (anonymous):

i got it -2,6

OpenStudy (jdoe0001):

i got it -2,6 \(\checkmark\)

OpenStudy (anonymous):

thanks